OpenStudy (anonymous):

how would you write these logarithmic expressions as a single logarithm : 1. 4 log m - log n 2. log 5 - k log 2 3.) log [ base of 6] 5 + log [ base of 6 ] x 4. log { base of 7 ] x + log [base of 7] y - log [base of 7] z

OpenStudy (zzr0ck3r):

a*ln(b) = ln(b^a) ln(a)-ln(b) = ln(a/b) ln(a)+ln(b) = ln(a*b) use these rules to get the first one and show me what you get

OpenStudy (anonymous):

i got log m^4 / n ..... is that right ?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

okay so how do i get the answer for the second one ?

OpenStudy (anonymous):

@Zekarias

OpenStudy (anonymous):

\[\log 5-\log 2^{k}\]\[\log(\frac{ 5 }{ 2^{k} })\]

OpenStudy (anonymous):

yea that is exactly wat i got bt now im at the third one and i dont know how to solve it

OpenStudy (anonymous):

\[\log_{6}5+\log_{6}x \]\[\log_{6}5x \]

OpenStudy (anonymous):

and then would the third one be : log[base 7] xyz

OpenStudy (anonymous):

?

OpenStudy (anonymous):

AND THE LAST ONE\[\log_{7} x+\log_{7}y-\log_{7}z\]\[\log_{7}(\frac{ xy }{ z })\]

OpenStudy (anonymous):

thank you so much for ur help